A solution of the diprotic acid hydrosulfuric acid is 0.10 M. Determine the equilibrium molarity of every species present. Can the overall reaction be used? Explain.
Answer #1

H2S(aq) -----------> H^+ (aq) + HS^- (aq)

I             0.1                           0                 0

C            -x                            +x                 +x

E           0.1-x                        +x                 +x

            Ka    =   [H^+][HS^-]/[H2S]

           9*10^-8   = x*x/(0.1-x)

           9*10^-8*(0.1-x) = x^2

            x   = 9.5*10^-5

           [H^+]   = x = 9.5*10^-5M

           [HS^-]   = x = 9.5*10^-5M

           HS^- (aq) ---------------> H^+ (aq) + S^2- (aq)

              Ka2    = [H^+][S^2-]/[HS^-]

              1*10^-17   = 9.5*10^-5[S^2-]/9.5*10^-5

             [S^2-]    = 1*10^-17M

              PH   = -log[H^+]

                     = -log(9.5*10^-5)

                      = 4.0222>>>>answer